OpenStudy (anonymous):

Suppose an asteroid with a diameter of 9.0km and a mass of 1.0×10^15kg hits the earth with an impact speed of 4.5×10^4m/s . What is the earth's recoil speed after such a collision? Assume that the mass of the earth is 5.98⋅10^24 kg. (Use a reference frame in which the earth was initially at rest.) What percentage is this of the earth's speed around the sun? Assume that the orbit of the earth is a circle with the radius of 1.5⋅10^11 m.

OpenStudy (irishboy123):

They say ""Use a reference frame in which the earth was initially at rest.)"" Thus it seems that you just need to distribute the momentum of the asteroid ma * va over it and the earth after collision ie (ma + me) v = ma*va, where v is what you are trying to find. for the second, distance of orbit = 2 π r metres time taken = 365 * 24 * 60 * 60 seconds. average speed = distance/time [a better way to do this is to compare the earth's angular moment about the sun before and after collision, and that could make the asteroid's density relevant.]

OpenStudy (anonymous):

According to Law of Conservation of Momentum Mass of asteroid*Speed of asteroid before impact = (Mass of Earth+Mass of asteroid)*speed after impact. 1. Hope you will get speed say v 2. Normal speed of Earth Vn = Distance traveled for one rotation/Time taken = 2*pi*r/one year compare between v and Vn Increase or decrease in speed of earth depends upon the direction of impact (according to Newtons second Law)